Todrick Company is a merchandiser that reported the following information based on 1,000 units sold: $ 210,000 $ 14,000 $ 140Req 1 Req 2 Req 3 to 5 Req 6 Prepare a contribution format income statement. Todrick Company Contribution Format Income StateReq 1 Req 2 Req 3 to 5 Req 6 Prepare a traditional format income statement. Todrick Company Traditional Income Statement SellReq 1 Req 2 Req 3 to 5 Req 6 3. Calculate the selling price per unit. 4. Calculate the variable cost per unit. 5. Calculate tReg 1 Req 2 Req 3 to 5 Req 6 Which income statement format (traditional format or con estimating how net operating income wil
Answer #1
1 Contribution format income statement
$ $
Sales 210000
Less: Variable expenses
Cost of goods sold (Note:1) 147000
Variable selling expenses 10500
Variable administrative expenses (Note:2) 10500 168000
Contribution margin 42000
Less: Fixed expenses
Fixed selling expenses (Note:3) 21000
Fixed administrative expenses 8400 29400
Net operating income 12600
Note:1
Cost of goods sold=Beginning merchandise inventory+Purchases-Ending merchandise inventory=14000+140000-7000=$ 147000
Note:2
Contribution margin=Sales-Variable expenses
Variable expenses=Sales-Contribution margin=210000-42000=$ 168000
Variable expenses=Sales-Cost of goods sold+Variable selling expenses+Variable administrative expenses
Variable administrative expenses=Variable expenses-Sales-Cost of goods sold-Variable selling expenses=168000-147000-10500=$ 10500
Note:3
Net operating income=Contribution margin-Fixed expenses
Fixed expenses=Contribution margin-Net operating income=42000-12600=$ 29400
Fixed expenses=Fixed selling expense+Fixed administrative expense
Fixed selling expense=Fixed expenses-Fixed administrative expense=29400-8400=$ 21000
2 Traditional format income statement
$ $
Sales 210000
Less: Cost of goods sold 147000
Gross profit 63000
Less:Selling and administrative expense
Variable selling expenses 10500
Variable administrative expenses (Note:2) 10500
Fixed selling expenses (Note:3) 21000
Fixed administrative expenses 8400 50400
Net operating income 12600
3 Selling price per unit=Sales/Units sold=210000/1000=$ 210
4 Variable cost per unit=Total variable expenses/Units sold=168000/1000=$ 168
5 Contribution margin per unit=Contribution margin/Units sold=42000/1000=$ 42
6 Contribution margin income statement